# Tony Lindeberg1    Zahra Babaiee2 and Peyman M. Kiasari2

### v1

@@ 1.1 Structure of this paper
In previous work, the use of such spatial spread measure has been demonstrated to have a predictive ability to compare the effective amount of spatial smoothing in receptive fields, when performing learning of the scale levels in Gaussian derivative networks using different types of discretization operations for Gaussian derivative operators, see Perzanowski and Lindeberg (2025).
@@ 2.1.1 Spatial spread measures applied to zero-order scale-space kernels
where $I$ denotes the unit matrix. Similarly, for the 2-D discrete analogue of the Gaussian kernel according to (Lindeberg 1990, 2024)
@@
where $I_{n}(\sigma^{2})$ denote the modified Bessel functions of integer order (see Abramowitz and Stegun 1964), the corresponding discrete spatial spread measure is correspondingly (Lindeberg 1993a, 2024)
@@
Since both the continuous Gaussian kernel and the discrete analogue of the Gaussian kernel are centered, we also have that the spatial mean values are
$\begin{split}M(g(\cdot,\cdot;\;\sigma))=0,\end{split}$ (11)
$\begin{split}M(T(\cdot,\cdot;\;\sigma))=0.\end{split}$ (12)
@@ 2.1.2 Spatial spread measures applied to first-order scale-space kernels
Later, we will make use these results in combination with computed discrete variances of discrete kernels, that resemble first-order derivative approximations, for estimating the scale parameter $\sigma$ in corresponding idealized models of receptive fields in terms of scale-space operations.
@@ 2.1.3 Spatial spread measures applied to zero-order pure difference operators
Filter $\|h_{i}(\cdot,\cdot)\|_{1}$ $M(|h_{i}(\cdot,\cdot))|$ $V(|h_{i}(\cdot,\cdot))|$
@@
Table 1: The result of computing the discrete $l_{1}$-norm $\|h(\cdot,\cdot)\|_{1}$ as well as the discrete spatial mean value $M(|h(\cdot,\cdot))|$ according to Equation (3) and the discrete spatial variance $V(|h(\cdot,\cdot))|$ according to Equation (4) for the 8 “master key filters” shown in Figure 1.
@@ 2.2 Measuring the spatial extent and the spatial offset of the 8 “master key filters”
as well as the discrete spatial spread measures $M(|h(\cdot,\cdot))|$ and $V(|h(\cdot,\cdot))|$ according to Equations (3) and (4) for the 8 “master key filters” extracted by Babaiee et al. (2025a) and shown in Figure 1. As can be seen from these results:
  • For all the filters, the off-diagonal elements in the spatial covariance matrices $V(|h(\cdot,\cdot))|$ are much smaller in magnitude than the diagonal elements. In this respect, the learned receptive field appear to be strongly aligned with the Cartesian coordinate directions.
  • For Filters 1–4, one component of the spatial mean value $M(|h(\cdot,\cdot))|$ is very close to zero, while the other component is rather close to either $\tfrac{1}{2}$ or $\tfrac{1}{2}$, thus reflecting a substantial deviation from symmetry. In this respect, there is a close structural similarity to the corresponding spatial mean values for the non-centered first-order difference operators according to Equations (28)–(31).
@@
  • For Filters 5–6 and Filters 7–8, the spatial mean values $M(|h(\cdot,\cdot))|$ are overall very close to zero, thus indicating a strong degree of either mirror symmetry or antisymmetry.
  • For all the Filters 1–8, the mixed component in the variance-based spatial spread measure $V(|h(\cdot,\cdot))|$ is very close to zero, indicating a very strong connection of the receptive field shapes to the preferred directions of the Cartesian coordinate system.
@@
  • For Filters 5–6, the ratio between the diagonal elements in the variance-based spatial spread measure $V(|h(\cdot,\cdot))|$ is very close to either 2 or $\tfrac{1}{2}$, indicating a rather close similarity to corresponding continuous variance-based spatial spread measures for first-order Gaussian derivative kernels according to Equations (15) and (18). The similarity is, however, not fully perfect, thus reflecting a minor notable difference.
@@
  • For Filters 1–4, the ratio between the diagonal elements in the variance-based spatial spread measure $V(|h(\cdot,\cdot))|$ is, however, not in any way close to either 2 or $\tfrac{1}{2}$, indicating that a direct transfer of the corresponding expressions (15) and (18) for continuous first-order Gaussian derivatives does not appear as straightforward.
@@
  • For Filter 8, the diagonal components of the variance-based spatial spread measure $V(|h(\cdot,\cdot))|$ are very similar, thus with a very good qualitative agreement with Equations (7) and (10).
@@
  • For Filter 7, the diagonal elements of the variance-based spatial spread measure $V(|h(\cdot,\cdot))|$ are somewhat similar, although differing in a clearly notable respect.
@@
Concerning the interpretation of in particular the variance-based spatial spread measures for Filters 7-8, it should, however, be noted that the definitions of these spatial spread measures assume a background magnitude value equal to zero, while Filters 7–8 in the set of “master key filters” have background magnitude levels that are significantly different from zero, thereby leading to substantial bias in the spatial spread measures, by substantially overestimating the amount of spatial extent of the filters. Thus, some modification is needed, to be able to apply the variance-based spatial spread measures $V(|h(\cdot,\cdot))|$ for estimating filter parameters in idealized models to be used for modelling Filters 7–8.
@@
A natural consistency requirement when constructing a centered discrete derivative approximation operator $\Delta_{x^{\alpha}y^{\beta}}$ that is to approximate the derivative operator $\partial_{x^{\alpha}y^{\beta}}$ of a given order $(m,n)\in{\mathbb{Z}}_{+}^{2}$ over a 2-D spatial domain is to require that the result of applying the discrete derivative approximation filter of order $(m,n)$ applied to a monomial of the same order leads to a response that is spatially constant, with specifically
@@
With regard to modelling the 8 “master key filters” in Figure 1, where we have according to the analysis in connection with Table 1 noted that the 8 “master key filters” (i) are not necessarily centered, and (ii) do not decrease as rapidly with increasing distance from the center as the discrete derivative approximation filters considered in Lindeberg (2024, 2025),  we do for the purpose of the forthcoming analysis in this paper generalize the criterion in Equation (41) into the shift-adjusted consistency requirement
$\Delta_{x^{\alpha}y^{\beta}}((x-m_{x})^{\alpha}\,(y-m_{y})^{\beta})=\alpha!\,\beta!,$ (43)
and generalize the criterion in Equation (42) into the shift-adjusted consistency requirement
@@
Specifically, these consistency requirements hold for all of the pure non-central or central difference operators $\delta_{x+}$, $\delta_{x-}$, $\delta_{y+}$, $\delta_{y-}$, $\delta_{x}$, $\delta_{y}$, $\delta_{xx}$, $\delta_{xy}$ and $\delta_{yy}$ that we consider in this work.
@@ 2.3.2 Responses to monomials for continuous Gaussian derivative operators
Due to the fact that the spatial scale-space representation of any image $f(x,y)$, obtained by Gaussian smoothing expressed in terms of a parameterization of the scale parameter in terms of the variance $s=\sigma^{2}$ of the Gaussian kernel according to
@@
if follows that evolution over scale for monomial input $f_{ab}(x,y)=x^{a}\,y^{b}$ can be described in terms of diffusion polynomials, where we for the lowest orders $a$ and $b$ have that
$\begin{split}q_{00}(x,y;\;s)=1,\end{split}$ (48)
$\begin{split}q_{10}(x,y;\;s)=x,\end{split}$ (49)
$\begin{split}q_{01}(x,y;\;s)=y,\end{split}$ (50)
$\begin{split}q_{20}(x,y;\;s)=x^{2}+s,\end{split}$ (51)
$\begin{split}q_{11}(x,y;\;s)=x\,y\end{split}$ (52)
$\begin{split}q_{02}(x,y;\;s)=y^{2}+s.\end{split}$ (53)
@@
$\begin{split}g_{y}(\cdot,\cdot)*y&=0.\end{split}$ (62)
@@
By further differentiating these relationsship, it follows that the first-order non-centered Gaussian derivative response in the $x$-direction to the monomial input $f_{10}(x,y)=x-m_{x}$ as well as the first-order non-centered Gaussian derivative response in the $x$-direction to the monomial input $f_{01}(x,y)=y-m_{y}$ will be equal to 1, if the centering for the non-centered Gaussian derivative responses is around the point $(m_{x},m_{y})^{T}$.
@@
1 $2.0\cdot 10^{-6} -$0.079 +1.556
2 $5.0\cdot 10^{-6} -$1.380 +0.396
3 $2.0\cdot 10^{-6}$ +0.129 -1.091
4 $5.6\cdot 10^{-18}$ +0.839 +0.029
5 $2.0\cdot 10^{-6}$ +7.750 -0.063
6 $4.0\cdot 10^{-6} -$0.160 +6.137
7 $1.0\cdot 10^{-6} +$0.085 -0.169
8 $4.0\cdot 10^{-16} -$0.072 -0.167
Table 2: The result of computing the responses to the monomials $\theta_{00}=1$, $\theta_{10}=x-m_{x}$ and $\theta_{01}=y-m_{y}$ for the 8 “master key filters” shown in Figure 1, where the spatial offset $(m_{x},m_{y})^{T}=M(|h(\cdot,\cdot)|$ represents the spatial mean of the absolute value of each filter.
@@
Table 3: The result of computing the responses to the monomials $\theta_{00}=1$, $\theta_{10}=x-m_{x}$ and $\theta_{01}=y-m_{y}$ for normalized versions of the first 6 “master key filters” shown in Figure 1, where the spatial offset $(m_{x},m_{y})^{T}=M(|h(\cdot,\cdot)|$ represents the spatial mean of the absolute value of each filter. For this table, the amplitudes of the original first 6 “master key filters” have been rescaled, to make the response equal to 1 for the monomial of $\theta_{10}=x-m_{x}$ or $\theta_{01}=y-m_{y}$ that best matches an idealized approximation of the filter as either a derivative in the $x$-direction or the $y$-direction.
@@ 2.3.3 Responses to monomials for the 8 “master key filters” operators
Table 2 shows the result of computing the responses to the monomials $\theta_{00}=1$, $\theta_{10}=x-m_{x}$ and $\theta_{01}=y-m_{y}$ for the 8 “master key filters” in Figure 1, with the spatial offset $(m_{x},m_{y})^{T}$ determined from the spatial mean of the absolute value of each filter according to $(m_{x},m_{y})^{T}=M(|h(\cdot,\cdot)|$. As can be seen from these results:
  • The responses of the filters to the constant function $\theta_{00}=1$ is very close to zero for all the filters, indicating that the filter are in this respect DC-balanced.
@@
  • For Filters 7 and 8, which have the qualitatively similar visual appearance as a discrete Gaussian kernel, or a discrete delta function with a subtracted Laplacian-of-the-Gaussian, the DC-balancing of the “master key filters” is, however, not fully consistent with discrete Gaussian models of those filters, since the both the continuous integral of the continuous Gaussian kernel and the discrete sum of the filter coefficients in discrete approximations of Gaussian kernels should be equal to 1
Table 3 shows the result of an extension of this analysis, where we have for the first 6 “master key filters” normalized the amplitudes of the receptive fields to make the response equal to 1 for the monomial of $\theta_{10}=x-m_{x}$ or $\theta_{01}=y-m_{y}$ that best matches an idealized approximation of the filter as either a derivative in the $x$-direction or the $y$-direction:
@@
    In this respect, the responses to first-order monomials is rather well consistent as approximations to first-order derivative operators for Filters 1, 4, 5 and 6.
@@
The matches to corresponding fully idealized models in terms of Gaussian derivatives or discrete approximations thereof is, however, far from perfect. A possible source to deviations from the idealized theory could also be because of from variations in the background periphery of the learned filters, which, by a lack of symmetry in those spurious variations, could contribute significantly to the responses to the first-order monomials.
@@
As previously remarked, for Filters 7 and 8, the estimates of the spatial extents of the filters will be significantly biased, because of the lack of a determination of the DC-level of these filters to correspond to a background level of about zero. Instead, as we saw from the experimental results in Table 2, the response of these filters to the constant function $\theta_{00}=1$ is very close to zero, as opposed to being close to one, as it should be for a continuous Gaussian filter or a discrete approximation thereof.
@@
for $i\in\{7,8\}$, and then further normalize the resulting DC-corrected filters according to
@@
Figure 5: Graphs of the error measure $\sqrt{\det V(|h_{i}(\cdot,\cdot)-C_{i}|)}$ for $i\in\{7,8\}$ when determining the DC compensation constants $C_{7}\approx-0.0118$ and $C_{8}\approx-0.0386$ according to (71), in order to later renormalize the Gaussian-like “master key filters” to unit $l_{1}$-norm, as regular Gaussian kernels obey.
Filter $\|h_{i}(\cdot,\cdot)\|_{1}$ $M(|h_{i,\scriptsize\mbox{DC}}(\cdot,\cdot))|$ $V(|h_{i,\scriptsize\mbox{DC}}(\cdot,\cdot))|$
@@
Table 4: The result of computing the discrete $l_{1}$-norm $\|h(\cdot,\cdot)\|_{1}$ as well as the discrete spatial mean value $M(|h(\cdot,\cdot))|$ according to Equation (3) and the discrete spatial variance $V(|h(\cdot,\cdot))|$ according to Equation (4) for the DC-compensated “master key filters” 7 and 8 according to (69) and (70).
@@
Table 5: The result of computing the responses to the monomials $\theta_{00}=1$, $\theta_{10}=x-m_{x}$ and $\theta_{01}=y-m_{y}$ for normalized versions of the last 2 “master key filters” shown in Figure 1, where the spatial offset $(m_{x},m_{y})^{T}=M(|h(\cdot,\cdot)|$ represents the spatial mean of the absolute value of each filter. For this table, the amplitudes of the original last 2 “master key filters” have been rescaled, to make the response equal to 1 for the monomial of $\theta_{00}=1$ equal to 1.
@@
according to (71) for determining the DC compensation constants for $\hat{C_{7}}$ and $\hat{C_{8}}$ in this way. Table 4 shows the result of computing the spatial spread measures for the resulting DC-compensated filters $h_{7,\scriptsize\mbox{DC}}(m,n)$ and $h_{8,\scriptsize\mbox{DC}}(m,n)$ according to (69) and (70). As can be seen from the results, the use of DC compensation substantially reduces the variance-based spatial spread measures for Filters 7 and 8, compared to the variance-based spread measures of the corresponding not DC-compensated filters shown in Table 1.
@@
Figure 2 shows a visualization of the resulting normalized versions of the 8 “master key filters”, which we will then approximate by idealized receptive field models in different ways in Section 3. Before proceeding to the modelling stage, we will, however, first in the next section introduce a way of defining spatial spread measures, that are less influenced by spurious variabilities in the peripheries of the learned filters.

### v2

@@ 2.4 Tangible contributions to ConvNeXt networks
The use of these spatial spread measures is structurally related to the use of mean values and covariance matrices for characterizing gross overall properties of statistical distributions. For our purpose, of capturing essential properties of both the learned filters in a deep network architecture and the use of idealized filter models to replace those filters, the motivation behind using these measures is to in a compact manner characterize how both the learned filters and the idealized filters weigh the image information from different spatial points.
In previous work, the use of such spatial spread measure has been demonstrated to have a predictive ability, to compare the effective amount of spatial smoothing in receptive fields, when performing learning of the scale levels in Gaussian derivative networks using different types of discretization operations for Gaussian derivative operators, see Perzanowski and Lindeberg (2025).
@@ 3.1.1 Spatial spread measures applied to zero-order scale-space kernels
where $I$ denotes the unit matrix. Similarly, for the 2-D discrete analogue of the Gaussian kernel (Lindeberg 1990, 2024)
@@
where $I_{n}(\sigma^{2})$ denotes the modified Bessel functions of integer order (see Abramowitz and Stegun 1964), the corresponding discrete spatial spread measure is correspondingly (Lindeberg 1993a, 2024)
@@
Since both the continuous Gaussian kernel and the discrete analogue of the Gaussian kernel are centered, we also have that their spatial mean values are
$\begin{split}M(g(\cdot,\cdot;\;\sigma))=\left(\begin{array}[]{c}0\\ 0\end{array}\right),\end{split}$ (11)
$\begin{split}M(T(\cdot,\cdot;\;\sigma))=\left(\begin{array}[]{c}0\\ 0\end{array}\right).\end{split}$ (12)
@@ 3.1.2 Spatial spread measures applied to first-order scale-space kernels
Later, we will make use of these results in combination with computed discrete variances of discrete kernels, that resemble first-order derivative approximations, for estimating the scale parameter $\sigma$ in corresponding idealized models of receptive fields in terms of scale-space operations.
@@ 3.1.3 Spatial spread measures applied to zero-order pure difference operators
Filter $\|h_{i}(\cdot,\cdot)\|_{1}$ $M(|h_{i}(\cdot,\cdot)|)$ $V(|h_{i}(\cdot,\cdot)|)$
@@
Table 1: The result of computing the discrete $l_{1}$-norm $\|h(\cdot,\cdot)\|_{1}$ as well as the discrete spatial mean value $M(|h(\cdot,\cdot)|)$ according to Equation (3) and the discrete spatial variance $V(|h(\cdot,\cdot)|)$ according to Equation (4) for the 8 “master key filters” shown in Figure 1.
@@ 3.2 Measuring the spatial extent and the spatial offset of the 8 “master key filters”
as well as the discrete spatial spread measures $M(|h(\cdot,\cdot)|)$ and $V(|h(\cdot,\cdot)|)$ according to Equations (3) and (4) for the 8 “master key filters” extracted by Babaiee et al. (2025a) and shown in Figure 1. As can be seen from these results:
  • For all the Filters 1–8, the off-diagonal mixed elements in the variance-based spatial spread measures in terms of the covariance matrices $V(|h(\cdot,\cdot)|)$ are much smaller in magnitude than the diagonal elements. In this respect, the learned receptive field appear to be strongly aligned with the Cartesian coordinate directions.
  • For Filters 1–4, one component of the spatial mean value $M(|h(\cdot,\cdot)|)$ is much smaller in magnitude than the other component, which is rather close to either $\tfrac{1}{2}$ or $\tfrac{1}{2}$, thus reflecting a substantial deviation from symmetry. In this respect, there is a close structural similarity to the corresponding spatial mean values for the non-centered first-order difference operators in Equations (28)–(31).
@@
  • For Filters 5–6 and Filters 7–8, the spatial mean values $M(|h(\cdot,\cdot)|)$ are rather close to zero, in the sense that their magnitudes are all smaller than 0.1, thus indicating a strong degree of either mirror symmetry or antisymmetry.
@@
  • For Filters 5–6, the ratio between the diagonal elements in the variance-based spatial spread measure $V(|h(\cdot,\cdot)|)$ is very close to either 2 or $\tfrac{1}{2}$, indicating a rather close similarity to corresponding continuous variance-based spatial spread measures for first-order Gaussian derivative kernels in Equations (15) and (18). The similarity is, however, not fully perfect, thus reflecting a minor notable difference.
@@
  • For Filters 1–4, the ratio between the diagonal elements in the variance-based spatial spread measure $V(|h(\cdot,\cdot)|)$ is, however, not in any way close to either 2 or $\tfrac{1}{2}$, indicating that a direct transfer of the corresponding expressions (15) and (18) for the continuous first-order Gaussian derivatives does not appear as straightforward.
@@
  • For Filter 8, the diagonal components of the variance-based spatial spread measure $V(|h(\cdot,\cdot)|)$ are very similar, thus with a very good qualitative agreement with Equations (7) and (10).
@@
  • For Filter 7, the diagonal elements of the variance-based spatial spread measure $V(|h(\cdot,\cdot)|)$ are somewhat similar, although differing in a clearly notable respect.
  • The discrete $l_{1}$-norms of Filters 1–4, which later will be shown to be different instances of the same class of filters, are of a similar order of magnitude. The discrete $l_{1}$-norms of Filters 5–6 are also of a similar magnitude. This property indicates that the data-driven way of determining the 8 “master key filters” can be regarded as reasonably robust between the clustered filters from the same class.
  • The discrete $l_{1}$-norms of all the filters $\{1,2,3,4,5,6,7,8\}$ are of a roughly similar order of magnitude, while differing between the different classes $\{1,2,3,4\}$, $\{5,6\}$, $\{7\}$ and $\{8\}$. [^5] While one could potentially consider normalizing the filters with respect to their $l_{1}$-norms, possibly considering the way that the $L_{1}$-norms of continuous Gaussian derivative kernels differ between different order of spatial differentiation, we will, however, not here normalize the filters with respect to their $l_{1}$-norms, and instead normalize the filters with respect to their responses to monomials, as will be later described in Section 3.3.3.
Concerning the interpretation of, in particular, the variance-based spatial spread measures for Filters 7-8, it should, however, be noted that the definitions of these spatial spread measures assume a background magnitude value equal to zero, while Filters 7–8 in the set of “master key filters” have background magnitude levels that are significantly different from zero. If we are to use the spatial spread measures for capturing the qualitative shapes of both learned and idealized filters in this way, a potential problem, however, is that the occurence of spurious non-zero values for the learned filters at image positions, where the their corresponding idealized filters are close to zero or very small, can lead to biases in the spatial spread measures. An explanation for this is that the contribution to the spatial spread measure will be weighted by the distance from the origin for the first-order spatial spread measure and by the distance squared for the second-order spatial spread measure.
@@
Thus, if we would match the parameters of the idealized filters to the shapes of the learned filters by regular spatial spread measures only, the resulting idealized filters could become spatially too large, thereby leading to substantial bias caused by the spatial spread measures, by substantially overestimating the spatial extent of the filters. Thus, some modification is needed, to be able to apply the variance-based spatial spread measures $V(|h(\cdot,\cdot)|)$ for estimating the filter parameters in the idealized models to be used for modelling Filters 7–8. We will return to this issue in Section 3.4 and Section 3.5.
@@
A natural consistency requirement, when constructing a centered discrete derivative approximation operator $\Delta_{x^{\alpha}y^{\beta}}$, that is to approximate the derivative operator $\partial_{x^{\alpha}y^{\beta}}$ of a given order $(m,n)\in{\mathbb{Z}}_{+}^{2}$ over a 2-D spatial domain, is to require that the result of applying the discrete derivative approximation filter of order $(m,n)$ applied to a monomial of the same order leads to a response that is spatially constant, with specifically
@@
With regard to modelling the 8 “master key filters” in Figure 1, we did according to the analysis in connection with Table 1 note that the 8 “master key filters” (i) are not necessarily centered, and (ii) do not decrease as rapidly with increasing distance from the center as the discrete derivative approximation filters considered in Lindeberg (2024, 2025). To handle the off-centering of the filters, we do for the purpose of the forthcoming analysis in this paper generalize the criterion in Equation (41) into the shift-adjusted consistency requirement
$\Delta_{x^{\alpha}y^{\beta}}((x-m_{x})^{\alpha}\,(y-m_{y})^{\beta})=\alpha!\,\beta!.$ (43)
We also generalize the criterion in Equation (42) into the shift-adjusted consistency requirement
@@
Specifically, these consistency requirements hold for all of the pure non-central or central difference operators, $\delta_{x+}$, $\delta_{x-}$, $\delta_{y+}$, $\delta_{y-}$, $\delta_{x}$, $\delta_{y}$, $\delta_{xx}$, $\delta_{xy}$ and $\delta_{yy}$, that we consider in this work.
@@ 3.3.2 Responses to monomials for continuous Gaussian derivative operators
Due to the fact that the spatial scale-space representation of any image $f(x,y)$, obtained by Gaussian smoothing expressed in terms of a parameterization of the scale parameter in terms of the variance $s=\sigma^{2}$ of the Gaussian kernel
@@
it follows that the evolution over scale for monomial input $f_{ab}(x,y)=x^{a}\,y^{b}$ can be described in terms of diffusion polynomials, where we for the lowest orders $a$ and $b$ have that
$\begin{split}f_{00}(x,y;\;s)=1,\end{split}$ (48)
$\begin{split}f_{10}(x,y;\;s)=x,\end{split}$ (49)
$\begin{split}f_{01}(x,y;\;s)=y,\end{split}$ (50)
$\begin{split}f_{20}(x,y;\;s)=x^{2}+s,\end{split}$ (51)
$\begin{split}f_{11}(x,y;\;s)=x\,y\end{split}$ (52)
$\begin{split}f_{02}(x,y;\;s)=y^{2}+s.\end{split}$ (53)
@@
$\begin{split}g_{y}(\cdot,\cdot)*y&=1.\end{split}$ (62)
@@
By further differentiating these relationsship, it follows that the first-order non-centered Gaussian derivative response in the $x$-direction to the monomial input $f_{10}(x,y)=x-m_{x}$ as well as the first-order non-centered Gaussian derivative response in the $y$-direction to the monomial input $f_{01}(x,y)=y-m_{y}$ will be equal to 1, if the centering for the non-centered Gaussian derivative responses is around the point $(m_{x},m_{y})^{T}$.
@@
1 $\cdot 10^{-5} -0.0$79 +1.556
2 $0\cdot 10^{-5}$ -1.380 +0.396
3 $0\cdot 10^{-5}$ +0.129 -1.091
4 $0\cdot 10^{-5}$ +0.839 +0.029
5 $0\cdot 10^{-5}$ +7.750 -0.063
6 $0\cdot 10^{-5}$ -0.160 +6.137
7 $\cdot 10^{-5} +0.0$85 -0.169
8 $\cdot 10^{-5} -0.0$72 -0.167
Table 2: The result of computing the responses to the monomials $\theta_{00}=1$, $\theta_{10}=x-m_{x}$ and $\theta_{01}=y-m_{y}$ for the 8 “master key filters” shown in Figure 1, where the spatial offset $(m_{x},m_{y})^{T}=M(|h(\cdot,\cdot)|)$ represents the spatial mean of the absolute value of each filter.
@@
Table 3: The result of computing the responses to the monomials $\theta_{00}=1$, $\theta_{10}=x-m_{x}$ and $\theta_{01}=y-m_{y}$ for normalized versions of the first 6 “master key filters” shown in Figure 1, where the spatial offset $(m_{x},m_{y})^{T}=M(|h(\cdot,\cdot)|)$ represents the spatial mean of the absolute value of each filter. For this table, the amplitudes of the original first 6 “master key filters” have been rescaled, to make the response equal to 1 for the monomial of $\theta_{10}=x-m_{x}$ or $\theta_{01}=y-m_{y}$ that best matches an idealized approximation of the filter as either a derivative in the $x$-direction or the $y$-direction.
@@ 3.3.3 Responses to monomials for the 8 “master key filters” operators
Table 2 shows the result of computing the responses to the monomials $\theta_{00}=1$, $\theta_{10}=x-m_{x}$ and $\theta_{01}=y-m_{y}$ for the 8 “master key filters” in Figure 1, with the spatial offset $(m_{x},m_{y})^{T}$ determined from the spatial mean of the absolute value of each filter according to $(m_{x},m_{y})^{T}=M(|h(\cdot,\cdot)|)$. As can be seen from these results:
  • The responses of the filters to the constant function $\theta_{00}=1$ are very close to zero for all the filters, indicating that the filters in this respect can be regarded as DC-balanced. [^6] In the greedy search procedure for extracting the “master key filters”, a normalization of the filter coefficients was used to having their sum equal to zero. In the data format used for storing the “master key filters” to be used as input for the modelling step, the filter coefficients were additionally rounded off to 6 decimals. Thereby, the responses of the filters to the monomial 1 will be quantized in units of $1\cdot 10^{-6}$. In Table 2, we have rounded off all these values to $0\cdot 10^{-5}$, implying consistency with the data-driven way that this analysis has been performed, without any use of such prior information, and the quantization errors and the floating-point errors caused by the finite precision in the storage of the “master key filters” and the computations on them.
@@
  • For Filters 7 and 8, which have the qualitatively similar visual appearance as a discrete Gaussian kernel, or a discrete delta function with a subtracted Laplacian-of-the-Gaussian, the DC-balancing of the “master key filters” is, however, not fully consistent with discrete Gaussian models of those filters, since both the continuous integral of the continuous Gaussian kernel and the discrete sum of the filter coefficients in discrete approximations of Gaussian kernels should be equal to 1
Table 3 shows the result of an extension of this analysis, where we have for the first 6 “master key filters” normalized the amplitudes of the receptive fields, to make the response equal to 1 for the monomial of $\theta_{10}=x-m_{x}$ or $\theta_{01}=y-m_{y}$ that best matches an idealized approximation of the filter as either a derivative in the $x$-direction or the $y$-direction:
@@
    In this respect, the responses to first-order monomials are rather well consistent as approximations to first-order derivative operators for Filters 1, 4, 5 and 6.
@@
The matches to the corresponding fully idealized models in terms of Gaussian derivatives or discrete approximations thereof is, however, far from perfect. A possible source to deviations from the idealized theory could also be because of from variations in the background periphery of the learned filters, which, by a lack of symmetry in those spurious variations, could contribute significantly to the responses to the first-order monomials.
@@
As previously remarked, for Filters 7 and 8, the estimates of the spatial extents of the filters obtained from direct application of the variance-based spatial spread measures will be significantly biased, because of the lack of a determination of the DC-level of these filters to correspond to a background level of about zero. Instead, as we saw from the experimental results in Table 2, the response of these filters to the constant function $\theta_{00}=1$ is very close to zero, as opposed to being close to one, as it should be for a continuous Gaussian filter or a discrete approximation thereof.
@@
for $i\in\{7,8\}$, and then further normalize the resulting DC-corrected filters as
@@
Figure 5: Graphs of the error measure $\sqrt{\det V(|h_{i}(\cdot,\cdot)-C_{i}|)}$ for $i\in\{7,8\}$ when determining the DC-compensation constants $C_{7}\approx-0.0118$ and $C_{8}\approx-0.0386$ according to (71), in order to later renormalize the Gaussian-like “master key filters” to unit $l_{1}$-norm, as regular Gaussian kernels obey. (Horizontal axes: parameter value $C_{i}\in[-0.1,0.1]$. Vertical axes: error measure: $\sqrt{\det V(|h_{i}(\cdot,\cdot)-C_{i}|)}$.)
Filter $\|h_{i,\scriptsize\mbox{DC}}(\cdot,\cdot)\|_{1}$ $M(|h_{i,\scriptsize\mbox{DC}}(\cdot,\cdot)|)$ $V(|h_{i,\scriptsize\mbox{DC}}(\cdot,\cdot)|)$
@@
Table 4: The result of computing the discrete $l_{1}$-norm $\|h(\cdot,\cdot)\|_{1}$ as well as the discrete spatial mean value $M(|h(\cdot,\cdot)|)$ according to Equation (3) and the discrete spatial variance $V(|h(\cdot,\cdot)|)$ according to Equation (4) for the DC-compensated “master key filters” 7 and 8 according to (69) and (70).
@@
Table 5: The result of computing the responses to the monomials $\theta_{00}=1$, $\theta_{10}=x-m_{x}$ and $\theta_{01}=y-m_{y}$ for normalized versions of the last 2 “master key filters” shown in Figure 1, where the spatial offset $(m_{x},m_{y})^{T}=M(|h(\cdot,\cdot)|)$ represents the spatial mean of the absolute value of each filter. For this table, the amplitudes of the original last 2 “master key filters” have been rescaled, to make the response equal to 1 for the monomial of $\theta_{00}=1$ equal to 1.
@@
according to (71) for determining the DC-compensation constants for $\hat{C_{7}}$ and $\hat{C_{8}}$ in this way. Table 4 shows the result of computing the spatial spread measures for the resulting DC-compensated filters $h_{7,\scriptsize\mbox{DC}}(m,n)$ and $h_{8,\scriptsize\mbox{DC}}(m,n)$ according to (69) and (70). As can be seen from the results, the use of DC compensation substantially reduces the variance-based spatial spread measures for Filters 7 and 8, compared to the variance-based spread measures of the corresponding not DC-compensated filters shown in Table 1.
@@
The top row in Figure 2 shows a visualization of the resulting normalized versions of the 8 “master key filters”, which we will then approximate by idealized receptive field models in different ways in Section 4. Before proceeding to the modelling stage, we will, however, first in the next section introduce a way of defining spatial spread measures, that are less influenced by spurious variabilities in the peripheries of the learned filters.